OpenStudy (anonymous):

Katie converts the outside temperature from degrees Fahrenheit, F, to degrees Celsius, C. She uses the formula below to convert the temperature. (F 32)5 9  C If the outside temperature is 50 degrees Fahrenheit, what is the outside temperature in degrees Celsius? F 2 G 5 H 9 J 10

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

(f-32) 5/9 = c is the formula so it turns into (50 - 32) 5/9 = C then (18) 5/9 = C

OpenStudy (skullpatrol):

yes, keep going :)

OpenStudy (texaschic101):

so whats 18 * 5/9 ?? Thats all you need, your almost done

OpenStudy (anonymous):

162 i think

OpenStudy (texaschic101):

whats 18 * 5 ? Then divide that by 9 and what do you get ?

OpenStudy (anonymous):

10

OpenStudy (texaschic101):

yep...you got it

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

you did it, not me. Way to go :)

OpenStudy (skullpatrol):

nice encouragement @texaschic101

OpenStudy (texaschic101):

thanks skullpatrol :)